OpenStudy (anonymous):

6.) Vera's piggybank contained 9.35 in quarters, dimes, and nickels. there were six more than two times as many nickels as quarters and four less dimes than quarters. how many of each kind of con was there in the bank?????? I don't need a answer. I need two equations..

jimthompson5910 (jim_thompson5910):

thanks

OpenStudy (anonymous):

your welcome

jimthompson5910 (jim_thompson5910):

let q = number of quarters d = number of dimes n = number of nickels

jimthompson5910 (jim_thompson5910):

we are told that "there were six more than two times as many nickels as quarters ", so we can say that q = 2n + 6

jimthompson5910 (jim_thompson5910):

there are " four less dimes than quarters.", so d = q - 4

jimthompson5910 (jim_thompson5910):

there is $9.35 in the bank, so 0.25q + 0.10d + 0.05n = 9.35 25q + 10d + 5n = 935 .. multiply everything by 100

jimthompson5910 (jim_thompson5910):

ok I made a typo, one sec

jimthompson5910 (jim_thompson5910):

The equation q = 2n+6 should be n = 2q+6

jimthompson5910 (jim_thompson5910):

25q + 10d + 5n = 935 25q + 10(q - 4) + 5n = 935 ... plug in d = q - 4 25q + 10q - 40 + 5n = 935 25q + 10q + 5n = 935+40 35q + 5n = 975 35q + 5(2q + 6) = 975 ... plug in n = 2q+6 35q + 10q + 30 = 975 45q + 30 = 975 45q = 975 - 30 45q = 945 q = 945/45 q = 21 So there are 21 quarters

jimthompson5910 (jim_thompson5910):

n = 2q+6 n = 2*21+6 n = 42 + 6 n = 48 so there are 48 nickels

jimthompson5910 (jim_thompson5910):

d = q - 4 d = 21 - 4 d = 17 and there are 17 dimes